I need the answer for the following or at least one example of how to do it
For the function, y= 1/x-2
Give the y values for x = -2, -1, 0, 1, 3.
Thanks

Answer by venugopalramana(3286)

For the function, y= 1/x-2 ...PUT BRACKETS ..I THINK YOU MEAN Y=1/(X-2) AND NOT (1/X)-2...
NOW SUBSTITUTE THE GIVEN VALUES OF X ONE BY ONE TO GET CORRESPONDING VALUES OF Y..SAY FOR X=0,WE GET
Y=1/(0-2)=1/(-2)=-1/2
FOR X=3,WE GET..
Y=1/(3-2)=1/1=1...ETC...
Give the y values for x = -2, -1, 0, 1, 3.
Thanks